Write an equation of the line perpendicular to the line given by 2x +3y =7 and having the same y-intercept as the line given 5x+ 2y =10
Answer by mananth(16949)

2x +3y =7 and having the same

y-intercept of line
5x+ 2y =10
2y=-5x+10
y=(-5/2)x +5
plug x=0
y=5
y intercept (0,5)
2 x + 3 y = 7
Find the slope of this line

3 y = -2 x + 7
Divide by 3
y = -0.667 x + 2.33
Compare this equation with y=mx+b
slope m = -0.67

The slope of a line perpendicular to the above line will be the negative reciprocal 1.5
m1*m2=-1
The slope of the required line will be 3/2

m= 3/2 ,point ( 0 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = 0.00 + b
b= 5
m= 3/2
The required equation isy = 3/ 2x + 5
m.ananth@hotmail.ca